OpenStudy (anonymous):

sharks teeth are shaped like triangles. the height of a tooth of a certain shark is 3 cm longer than the base. the area is 44 cm^2. Find the height and base.

OpenStudy (anonymous):

Area = \[\frac{ 1 }{ 2 }(base)(height)\] so we know the height is 3cm longer than the base, in other words: \[height = (base + 3cm)\] and so \[44cm ^{2} = \frac{ 1 }{ 2 }(base)(base + 3cm)\] find out the base, then plug back the value of the base into the original equation to find the height.

OpenStudy (wolf1728):

A) height = 3 cm + base B) .5 * h * b = 44 h = 44/(.5*b) h = 88/b multiplying this by -1 -h = -88/b then adding this to A) h = 3 + b 88/b = 3 + b multiplying both sides by b 88 = 3b + b² b² + 3b - 88=0 b= -3 +-sqrt(9 - 4*1-88)/2 b=8 h=11

OpenStudy (anonymous):

Thank you! :)